A six pack of soda costs $2.53. A carton of 24 cans costs $7.25. How much will be saved by purchasing the carton rather than 4 individual six packs?
Answer by checkley77(12844)

4*2.53=10.12-7.28=$2.84 is the savings when buying the carton.